Helen’s Health Foods usually sells 400 cans of ProPac Muscle Punch per week when the price is $5 per can. After experimenting with prices for some time, Helen has determined that the weekly demand can be found using the equation
d = 600 – 40 p
where d is the number of cans and p is the price per can.
a.) Will Helen sell more or less Muscle Punch if she raises her price from $5.00?
b.) What happens to her sales every time she raises her price by $1.00?
c.) Graph the equation.

Answer by Mathtut(3670)

a. d=600-40(5)=400 d=600-40(5.50)=380
as the price increases the the # of cans decrease
b. sales at $5 is 400(5)=2000
sales at $6 is 360(6)=2160
sales at $7 is 320(7)=2240
sales at $7.50 300(7.5)=2250
sales at $8 is 280(8)=2240

so sales rise on each $1 raise up to $7.50 then they begin to fall again.
c.I have no way to graph this